OpenStudy (anonymous):

Find the number of moles in 85 g KCl. A.3.3 mol KCl B.1.1 mol KCl C.11 mol KCl D.2.5 mol KCl

OpenStudy (anonymous):

1 mole of KCl has 39.10g for K and 35.45g for Cl Add these together: 74.55g Now 85g of KCl x 1 mol KCl / 74.55g = 1.1 mol Answer is B

OpenStudy (anonymous):

Well I took K as 38 and Cl as 34... So the relative molecular mass is.. \[R.m.m={(38*1)+(34*1)}\]\[=38+34\]\[=72\]So molar mass = 72gmol^-1 We could apply this in the formula\[n=\frac{m}{M}\]Where n = number of moles , m = molar mass of substance and M = mass of substance..So\[n=\frac{85g}{72gmol^{-1}}\]\[n=1.18mol\]\[n=1.1mol\].. :D